OpenStudy (anonymous):

Graph the exponential function. y=3(2)^x+4

OpenStudy (anonymous):

https://www.desmos.com/calculator could this help you?? If not sorry!

OpenStudy (anonymous):

its a start, thank you~

OpenStudy (whpalmer4):

\[y=3(2)^x + 4\] I would make a table of x and y values before breaking out the graph paper: x y -5 3(2)^-5 + 4 = 3/32 + 4 = 4.09375 -4 3(2)^-4 + 4 = 3/16+4 = 4.1875 -3 3(2)^-3 + 4 = 3/8 + 4 = 4.375 -2 3(2)^-2 + 4 = 3/4 + 4 = 4.75 -1 3(2)^-1 + 4 = 3/2 + 4 = 5.5 0 3(2)^0 + 4 = 3 + 4 = 7 1 3(2)^1 + 4 = 6+4 = 10 2 3(2)^2 + 4 = 12 + 4 = 16 3 3(2)^3 + 4 = 24 + 4 = 28 4 3(2)^4 + 4 = 48 + 4 = 52 5 3(2)^5 + 4 = 96 + 4 = 100 and that's probably more than sufficient. Now sharpen the pencil and start plotting points!

OpenStudy (whpalmer4):

As the x value goes toward - infinity, the y value gets closer and closer to 4, but never quite reaches it. After the x value passes 0 going in a positive direction, the function gets big really quickly :-)

OpenStudy (anonymous):

WHOA Thanks whpalmer4!!!

OpenStudy (whpalmer4):

Here's a graph (I like making graphs! :-) Don't look until you've plotted yours!

OpenStudy (anonymous):

I won't look, I promise~ :)

OpenStudy (anonymous):

i'm plotting and I think I'm going above the horizontal asymptote?

OpenStudy (anonymous):

oh I'm supposed to, oops I see why now